If the rate is at 1.25% effective, how long will R5 000 become R5 500?

1
Expert's answer
2021-12-07T11:22:13-0500

effective interest rate:

"r=(1+i\/n)^n-1"

where i is nominal rate,

n is number of compounding periods per year


we have:

"5500=5000(1+i\/n)^{nm}"

where m is number of years


"(1+i\/n)^n=r+1=1.0125"

then:

"5500=5000\cdot1.0125^m"


"m=ln1.1\/ln1.0125=7.67" years